Find answers to the following questions:
1. C(6,2)/5!
2.5!/2!*0!
3. 8!/2!*4!
4. 7!/P

Answer by Boreal(15235)

1. 6C2/5!=15/120=1/8
2. 5!/2!*0!= 120/2=60; 0!=1
3..8!/(2!4!)=40320/48=840. if it is( 8!/2!)*4!, it is 20160*24=483840. Parentheses matter
4. 7!/P=5040/P